# Sharp propagation of chaos in Rényi divergence

### v1

@@ 1 Introduction
$\mathcal{E}(\mu)=\int V\,\mathrm{d}\mu+\iint W(x-y)\,\mathrm{d}\mu^{\otimes 2}(x,y)+\int\log\mu\,\mathrm{d}\mu\,.$ (1.1)
@@
If $\mu$, $\pi$ satisfy sufficiently strong isoperimetric and smoothness conditions, then $\mathsf{R}_{q}(\mu^{1}\mathbin{\|}\pi)\lesssim\widetilde{O}(\frac{dq^{2}}{N^{2}})$, for $N$ sufficiently large.
@@ 1.1 Related work
The study of propagation of chaos has a venerable history which is not easily summarized; we touch only upon the most relevant works below. Beginning with the seminal work of Sznitman (2006) and Malrieu (2001, 2003), quantitative chaos bounds on the order of $1/N$ in $W_{2}^{2}$ or $\operatorname{\mathsf{KL}}$ divergence have been shown by first bounding the divergence between $\mu^{1:N}$ and $\pi^{\otimes N}$ by something that is independent of $N$, and then applying subadditivity. This line of thought has been successfully generalized to the case of singular interactions of $1/\lVert x^{i}-x^{j}\rVert^{\alpha}$ type (Jabin and Wang, 2017, 2018; Bresch et al., 2023), $\alpha>0$, which permits the study of Coulomb and Riesz gases, as well as other physical systems. See Chaintron and Diez (2022) for a more complete survey.
@@
The most relevant work to ours is Hess-Childs and Rowan (2025), which also provides a means of establishing higher order propagation of chaos in the dynamical setting. When the interaction is a.s. bounded, they show the sharp rate of $1/N^{2}$ via hierarchical techniques reminiscent of Lacker (2023) in the $\chi^{2}$-divergence (equivalent to $2$-Rényi), and which do not contain a “weak interaction” assumption. However, their analysis cannot be easily adapted to the case of unbounded-but-smooth interactions considered in this work. Finally, Bresch et al. (2022) offer a different proof for dynamical propagation of chaos in some $\chi^{p}$-norms, but which only works for short times and without quantitative $N$-dependence.
@@ 2.3 Main results
Remark. This requires that the finite-particle measure $\mu^{[k]}$ satisfy ($\mathsf{LSI}$) with a constant independent of $N$, for which it suffices that $\mu^{[N]}$ satisfy ($\mathsf{LSI}$) with the same constant. See §3 for some discussion of when this holds.
@@
We recall a standard change of measure lemma: for any event $A$ and two measures $\mu,\nu$, with $q>1$,
@@
We can recover this result by first applying exponential concentration results for the empirical measure under $\pi^{\otimes k}$, using techniques from Fournier and Guillin (2015), and then using Corollary 2 to guarantee, for $N$ large enough, the same result for the empirical measure under $\mu^{[k]}$. The final result is, up to the constant prefactor and the constant in the exponent, the same behaviour as verified in the setting of Jackson and Zitridis (2024). Although such a result no longer holds when $k=N$, it is an interesting illustration of the utility of $\mathsf{R}_{q}$ bounds.
@@ 3 Examples
Then the stationary distributions are Gaussians, and we can compute $\mathsf{R}_{q}(\mu^{1:k}\mathbin{\|}\pi^{\otimes k})=\widetilde{\Omega}(\tfrac{dk^{2}q}{N^{2}})$ for $N$ sufficiently large, which shows that our upper bound is tight with respect to the $N$ and $d$ scaling, assuming we keep $k,q$ relatively constant.
The same calculations as Kook et al. (2024, Example 10) give, for $\mathsf{C}\in\mathbb{R}^{N\times N}$ with $C_{i,i}=N-1$ and $C_{i,j}=-1$ if $i\neq j$, or in other words $\mathsf{C}=NI_{N}-J$, with $J=\operatorname{\mathbbm{1}}_{N}\!\operatorname{\mathbbm{1}}_{N}^{\top}$, $\operatorname{\mathbbm{1}}_{k}$ being the length $k$ vector with all entries equal to $1$,

### v2

@@ 1 Introduction
$\mathcal{E}(\mu)=\int V\,\mathrm{d}\mu+\frac{1}{2}\iint W(x-y)\,\mathrm{d}\mu^{\otimes 2}(x,y)+\int\log\mu\,\mathrm{d}\mu\,.$ (1.1)
@@
If $\mu$, $\pi$ satisfy sufficiently strong isoperimetric and smoothness conditions, then $\mathsf{R}_{q}(\mu^{1}\mathbin{\|}\pi)\lesssim\widetilde{O}(\frac{dq^{2}}{N^{2}})$, for NN sufficiently large.
@@ 1.1 Related work
The study of propagation of chaos has a venerable history which is not easily summarized; we touch only upon the most relevant works below. Beginning with the seminal work of Sznitman (2006) and Malrieu (2001, 2003), quantitative chaos bounds on the order of $1/N$ in $W_{2}^{2}$ or $\operatorname{\mathsf{KL}}$ divergence have been shown by first bounding the divergence between $\mu^{1:N}$ and $\pi^{\otimes N}$ by something that is independent of NN, and then applying subadditivity. This line of thought has been successfully generalized to the case of singular interactions of $1/\lVert x^{i}-x^{j}\rVert^{\alpha}$ type (Jabin and Wang, 2017, 2018; Bresch et al., 2023), $\alpha>0$, which permits the study of Coulomb and Riesz gases, as well as other physical systems. See Chaintron and Diez (2022) for a more complete survey.
@@
The most relevant work to ours is Hess-Childs and Rowan (2025), which also provides a means of establishing higher order propagation of chaos in the dynamical setting. When the interaction is a.s. bounded, they show the sharp rate of $1/N^{2}$ via hierarchical techniques reminiscent of Lacker (2023) in the $\chi^{2}$-divergence (equivalent to 22-Rényi), and which do not contain a “weak interaction” assumption. However, their analysis cannot be easily adapted to the case of unbounded-but-smooth interactions considered in this work. Finally, Bresch et al. (2022) offer a different proof for dynamical propagation of chaos in some $\chi^{p}$-norms, but which only works for short times and without quantitative NN-dependence.
@@ 2.3 Main results
Remark. This requires that the finite-particle measure $\mu^{[k]}$ satisfy ($\mathsf{LSI}$) with a constant independent of NN, for which it suffices that $\mu^{[N]}$ satisfy ($\mathsf{LSI}$) with the same constant. See §3 for some discussion of when this holds.
@@
We recall a standard change of measure lemma: for any event AA and two measures $\mu,\nu$, with $q>1$,
@@
We can recover this result by first applying exponential concentration results for the empirical measure under $\pi^{\otimes k}$, using techniques from Fournier and Guillin (2015), and then using Corollary 2 to guarantee, for NN large enough, the same result for the empirical measure under $\mu^{[k]}$. The final result is, up to the constant prefactor and the constant in the exponent, the same behaviour as verified in the setting of Jackson and Zitridis (2024). Although such a result no longer holds when $k=N$, it is an interesting illustration of the utility of $\mathsf{R}_{q}$ bounds.
@@ 3 Examples
Then the stationary distributions are Gaussians, and we can compute $\mathsf{R}_{q}(\mu^{1:k}\mathbin{\|}\pi^{\otimes k})=\widetilde{\Omega}(\tfrac{dk^{2}q}{N^{2}})$ for NN sufficiently large, which shows that our upper bound is tight with respect to the NN and $d$ scaling, assuming we keep $k,q$ relatively constant.
The same calculations as Kook et al. (2024, Example 10) give, for $\mathsf{C}\in\mathbb{R}^{N\times N}$ with $C_{i,i}=N-1$ and $C_{i,j}=-1$ if $i\neq j$, or in other words $\mathsf{C}=NI_{N}-J$, with $J=\operatorname{\mathbbm{1}}_{N}\!\operatorname{\mathbbm{1}}_{N}^{\top}$, $\operatorname{\mathbbm{1}}_{k}$ being the length $k$ vector with all entries equal to 11,